This year, Lambert Company will ship 1,500,000 pounds of goods to customers at a cost of $1,200,000. If a customer orders 10,000 pounds and produces $200,000 of revenue (total revenue is $20 million), the amount of shipping cost assigned to the customer by using ABC would be

a. unable to be determined. b. $8,000 ($0.80 per pound shipped). c. $24,000 (2% of the shipping cost). d. $12,000 (1% of the shipping cost). e. None of these.
Answer #1

Shipping cost assigned to customer

shipping cost per pound = $1200000/1500000=0.8 per pound

No of pound shipped= 10000

Shipping cost = 0.8*10000=$8000

Answer b.

In activity based costing Overhead allocate on the basis of cost driver .In this case cost driver is no of pounds shipped.